Safety regulations require that the time between airplane takeoffs (on the same runway) will be at least 3 minutes. When taking off, the run time of an airplane on the runway is 45 seconds. Planes are on average waiting 4 minutes and 15 seconds for take-off. On average there are 15 planes taking off per hour. How many planes are either on the runway or waiting to take off? (PLEASE SHOW WORK)
Answer #1

TR = 15 plans per hour.
TT = (0.75 +4.25)/60 = 1/12 of an hour
WIP = TT*TR = 15*1/12 = 1.25